• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name Mina Kat

    Log entry time 00:18:00 on April 9,2007

    Entry number 198291

    keyword=Shift Summary

    16:08 Javier, Ed in Hall setting Left spectrometer to desired angle
    of 139.52 for Q2 = 40fm-2 He3 running
    16:17 Done at the Hall. Called MCC asked for harp scans.

    16:25 Instead of being in the no target (empty position), the target ladder seems to have stuck in a position higher up and closer to the last solid target (copper).
    Requested from MCC to mask target movement. Moved target to "Home" position.
    Target "Home" position seems to be the empty target position
    (bottom of target ladder), but copper and carbon positions are in between targets.

    Put BeO target in.

    16:55 asked MCC to do harp scans.

    HVs were checked and trigger dowload was checked and found ok.
    Copies of the HV screens were put in the "E04-018 records" black binder.

    17:27 Harp scans finished and checked by us. Sigma x and sigma y are
    ok. Asked to insert OTR 1H04A.
    17:40 No success with OTR, MCC Crew Chief says that OTR is not in,
    it was removed Apr. 2nd.
    17:50 Decided to request beam at 20 uA to start Q2 = 40fm-2 data taking.

    Runs 2211 and 2212 short test runs. Spot++ ok

    Q2 = 40fm-2 productions runs:

    Run 2213 good run, current increased from 60 to 84 uA (during run)
    Runs 2213-2223 production runs
    22:05 Noticed 100 pct dead time, TS0 crash
    during run 2224 so no end run.
    22:05 Restarted CODA
    Run 2225 short test run
    Runs 2226-2231 production runs